OpenStudy (anonymous):

use a system of equations in two variables solve the following problem: 4 times a first number decreased by 3 times a second number is 22. The sum between the first number and twice the second number is 0. Find the number?

OpenStudy (anonymous):

first number = x second number = y so you have the equations: 4x - 3y = 22 x + 2y = 0

OpenStudy (anonymous):

Which number is it asking you to find? The first or the second?

OpenStudy (anonymous):

I have to get numbers for both

OpenStudy (anonymous):

Ok then lets solve for x first: 4x - 3y = 22 +3y 4x = 22 + 3y /4 x = 5.5 + 0.75y Now substitute the value of x into the second equation: x + 2y = 0 (5.5 + 0.75y) + 2y = 0 Now you combine the like terms: 5.5 + 0.75y + 2y = 0

OpenStudy (anonymous):

When you combine 5.5 + 0.75y + 2y = 0 you'll get the value of y.

OpenStudy (anonymous):

So what do you get when you solve fore y in: 5.5 + 0.75y + 2y = 0

OpenStudy (anonymous):

I get -2

OpenStudy (anonymous):

and for x i get 4 is this all correct?

OpenStudy (anonymous):

Thanks you if so.

OpenStudy (anonymous):

Yep your're right!:D No problem:)